# Polarimetric compressed sensing with hollow, self-assembled diffractive films

### v1

@@ Abstract
Sensing light’s polarization and wavefront direction enables surface curvature assessment, material identification, shadow differentiation, and improved image quality in turbid environments. Traditional methods utilize multiple sensors and polarization-filtering optics, resulting in reduced image resolution due to the loss of signal information at each photodetector. We propose a compressive sensing technique that mitigates information loss by using a low-refractive-index, self-assembled optical encoders. These optical nanostructures scatter light into polarization-dependent lattice modes and respond to input polarization ellipticity. Our study reveals that the wavefront direction and the polarization are encoded in the interference patterns and the linear component of the diffraction. Combining optical encoders with a shallow neural network, the system predicts pointing and polarization when equally spaced sensors are separated by less than 20°. Comparison of polycrystalline and short-range ordered optical encoders shows that the latter has a higher compression ratio. Our work centers on the unexpected modulation and spatial multiplexing of incident light polarization by self-assembled hollow nanocavity arrays as a new class of materials for optical computing, compressed sensing and imaging.
@@
American Chemical Society,
@@ 1 Introduction
Sensitivity to both light polarization and its wavefront or pointing direction offers the ability to infer the surface curvature, material, or texture, distinguish shadows from objects, see otherwise-transparent shapes, or improve image quality in turbid environments 1, 2, 3, 4. However, as independent qualities of light, polarization and pointing are typically detected separately with arrays of sensors and polarization-filtering optics. Polarization states are characterized by the Jones or Stokes parameters, i.e., four separate readings that identify the power (${\rm S}_{0}$), 2-$D$ linear-polarization (LP) axis (${\rm S}_{1}$, ${\rm S}_{2}$), and circular-polarization (CP) ellipticity (${\rm S}_{3}$). Since signal information is lost at each polarization-filtered photodetector, polarization-sensitive cameras involve inevitable reductions in image resolution.
The process of physical encoding would mitigate the challenge of lost information caused by polarization sensitivity through compressive sensing. Compressive sensing is a revolutionary technique that exploits the sparsity of a signal so that the data acquisition and storage requirements are significantly reduced without sacrificing the reconstruction quality 5. By contrast, the relatively cumbersome direct-sampling approach [Fig. 1(a)] involves rasterized or serial measurements to identify pointing and polarization with polarization filters and multiple-pixel sensors. A physical encoder that aids compressed sensing overlaps the input optical fields over wide areas of the output fields in a manner analogous to computing algorithms.
@@
where the sensing matrix $\mathbf{A}$ transfers the vector of input measurements $\mathbf{x}$ to a $1\times 4$ sensor output $\mathbf{y}$: this sensing approach is inefficient because many measurements are needed and there are many zero-valued elements in the sensing matrix $\mathbf{A}$. These zero-valued elements signify that the algorithmic operations performed yield no information.
Figure 1: Basic schematic of conventional polarimetric imaging and proposed polarimetric compressed sensing. (a) With a conventional polarimetric camera, the direction and polarization of a point source are measured with independent and separate filters and pixels. Each image vector $\mathbf{x}$ contains many zero-valued readings. (b) With encoded and hybrid sensing, the point light source and its polarization is mapped to multiple pixels with different polarizations. Each encoded image vector $\mathbf{\alpha}$ carries joint statistics of the pointing and polarization.
Figure 2: Characterization of the optical encoders. Digital image (a) and SEM image (b) of the PANI-HS optical encoder. (c) Cartoon illustration showing the cross section of the PANI-HS optical encoder. (d) Microscope image analysis of ‘ordered’ (top) and ‘random’ samples (bottom) showing the grain axes rotation (degrees). (e) The pair correlation function $g(r)$ computed from optical images against the normalized radius $r/2r_{0}$, where $r$ is the distance from an arbitrary origin, and $r_{0}$ is the mean sphere radius. The full width at half maximum (FWHM) of the first peak in the Fourier transform of the function $g(r)^{-1}$ provides a measure of the structural disorder. Perfect lattice (top), example ‘ordered’ (middle) and ‘random’ sample (bottom) are shown.
@@
Here, we demonstrate the compressed sensing of polarized signals with low-refractive-index, self-assembled encoders. The unexpected transformations from linear to elliptical polarizations in the colloidal grating structures have only recently been reported in our work 7. We draw an analogy to the compound-lens imaging system of insects, which is covered with self-assembled corneal random and ordered grating nanostructures 8, 9, 10. Although nanostructures in nature are associated with anti-reflective, hydrophobic, or self-cleaning functions 11, 12, the optical encoding function of their nanostructures has not been explicitly shown. A survey of the potential modes for insect-inspired imaging and object tracking (colorimetric, polarimetric filtering, coherent optical preprocessing, and compressive sensing) has recently been written 13. Whether insect corneal nanostructures have a physical encoding function or not, our approach with self-assembled encoders illustrates the potential.
A key component to our encoders is that they scatter light into polarization-dependent lattice modes and are sensitive to the input polarization ellipticity. Our work not only expands on our understanding of the physical encoding that is available with self-assembled nanostructures, but also outlines general insight for their design for compressed sensing.
@@ 2 Results
Figure 3: Polarization transformations of PANI-HS compared with those of polystyrene spheres (PS). (a) (i-ii) SEM images of PS and PANI-HS (720-nm lattice spacing) and corresponding (iii-iv) Debye or diffracted ring intensity patterns from vertical, linearly polarized (LP) light (532-nm wavelength). The measured diffracted polarization is overlaid in yellow. For PS, the diffracted intensity is uniform and the diffracted polarization remains vertical but for PANI-HS, both intensity and polarization vary spatially and the polarization ellipticity indicates spatially varying birefringence. (b) Conoscopy measurements showing (i-ii) diattenuation and (iii-iv) phase retardance magnitude and phase for single-crystalline PANI-HS, polycrystalline PANI-HS, and PS with 2-$\mu$m lattice spacing. Phase retardance and diattenuation are negligible in the PS samples compared to those of PANI-HS.
@@
We fabricated hierarchical 2D photonic-crystal films composed of polyaniline nanofibers around hollow spheres (PANI-HS) based on an approach demonstrated in our prior work 7. Fig. 2(a-c) show the air-filled polymer grating structures. The film structure at the smallest level is composed of nanofibers surrounding a sub-micron-scaled spherical cavity, as evidenced by the SEM image in Fig. 2(b). We prepared PANI-HS films with two types of structures: polycrystalline (’ordered’) and short-range ordered (’random’), to investigate the impact of ordering on compressive sensing. Structural analysis was conducted using boundary counts and the 2D pair correlation function. In the boundary count method 14, a grain is defined within a lattice with specific orientation; for our 6-fold symmetric assembly, a grain’s orientation relative to an axis ranges from 0 to 60 degrees. A grain boundary occurs when neighboring domains deviate by a small angle. Fig. 2(d) shows the orientation of different grains in a 30 $\mu$m × 22 $\mu$m region, revealing boundary counts of 1.8 boundaries/$\mu$m and 3.3 boundaries/$\mu$m for the ordered and random samples, respectively. Fig. 2(e) shows the 2D pair correlation function $g(r)$ 15, 16, 17 of the ordered and random samples alongside a perfect lattice. We assessed ordering by comparing the width of the first-order peaks in the Fourier transform to that of a perfect lattice, yielding average orderings of 1.10 and 2.49 for ordered and random samples across three randomly selected areas. Additional detailed structural analysis is provided in Supporting Information Fabrication and Material Characterization of PANI-HS (Fig. S1 and S2).
When light illuminates a larger, polycrystalline encoder area, we observe a diffracted ring also referred to as a Debye ring. The Debye ring and diffuse background patterns contain mixed, spatially multiplexed information, which arise from the mesostructure and the presence of nanofibers of the film. Figure 3 illustrates the polarization-sensitive Debye ring from PANI-HS, which enables us to differentiate RCP, LCP, and unpolarized light with an LP camera. This polarization-modulated diffraction does not occur with traditional colloidal structures; while the diffracted light from polystyrene spheres (PS) colloids (and most solid colloidal structures) is ring-shaped with unchanged polarization, the PANI-HS Debye ring intensity is broken and its polarization is inhomogeneous [Fig. 3(a)]. The hollow-sphere structure is important 18, 19: in fact, prior to the removal of PS, the PANI-coated, PS films exhibit trends that follow that of PS rather than PANI-HS. PS alone exhibits higher intensities in the Debye ring compared to PANI but does not transform the diffracted light polarization significantly. The changes in linear to elliptical polarizations are unexpected with colloidal grating structures. In fact, the polarization modulated diffraction is, to our knowledge, only recently reported in our work 7.
Figure 4: Modeled and measured polarizations a normal incidence. (a) Model description: the fields at each location on the diffracted Debye ring are a sample of the input beam filtered with a linear polarizer and wave plate. Both the polarizer and the slow-axis of the wave plate are aligned with the $k$-vector of the diffracted fields. Our model fits a 1:3 ratio between the polarizer and wave plate, and a phase retardance and relative phase associated with the diffractive resonance. (b) Stokes parameters of incidence changing from -90 (TE) to 0 (TM), back to 90 (TE). (c) Modeled and measured polarization transformations at the Debye ring angle (56).
@@
Angle-dependent diattenuation (or polarization-sensitive losses) and phase retardance are confirmed with Mueller matrix diffractometry 20 [Fig. 3(b)]. In these measurements, the Mueller matrix polarimeter images the rear focal plane of the 50x (0.95 NA) objective to obtain complete polarization-resolved maps of the directional distribution of light. The complete Mueller matrix for each image in Fig. 3(b) is provided in Supporting Information Diffractometry Measurements (Fig. S3). The comparison of measurements that probe single crystals vs. grain boundaries across multiple crystals [Figs. 3(b)i-ii] shows similar diattenuation and phase retardance. This symmetry of the diattenuation and retardance allows us to rule out stress-induced birefringence in the coating 21. Also, since similar trends are observed for single and multiple crystals, the birefringence is not associated with the grain boundaries or defects in the self-assembled films 22.
@@
The diffraction from nanofibrous PANI-HS films appears as the “pooled” scattered light from a single dipole scattering event with additional birefringence. A simple, coherent model fit to our measurements identifies scattering and phase retardance as the two underlying contributions to the polarization transformation. The polarization transformation is viewed as a rotated polarizer $\mathbf{J}_{\rm pol}$ and rotated waveplate $\mathbf{J}_{\rm wp}$,
@@
where $\phi_{0}$ is measured in the sample plane with respect to the $x$-axis, and $\mathbf{R}(\phi_{0})=\begin{bmatrix}\cos(\phi_{0})&-\sin(\phi_{0})\\ \sin(\phi_{0})&\cos(\phi_{0})\end{bmatrix}$ is the rotation matrix associated with azimuthal location $\phi_{0}$.
We convert our model from Jones to Stokes parameters. The input and diffracted Jones vectors $\mathbf{E}_{in}=E_{o}[A_{x},A_{y}\exp(-i\delta)]^{T}$ are shown with Stokes measures $S_{1}=A_{x}^{2}-A_{y}^{2}$, $S_{2}=2A_{x}A_{y}\cos(\delta)$, and $S_{3}=2A_{x}A_{y}\sin(\delta)$, where $A_{x}^{2}+A_{y}^{2}=1$. Additionally, $\delta=\pm\pi/2$ and $S_{3}\pm 1$ for RCP/LCP, respectively.
Subsequently, we interrogate the PANI-HS sample with linearly polarized light at normal incidence. The experimental Stokes parameters are measured by the polarimeter at a fixed location as the input polarization is changed with a half-wave plate. Figure 4(a) shows the experimental geometries and the corresponding Stokes parameters corresponding to the input beam (Figure 4(b)). The graph shows the transmitted beam reference Stokes parameters without the sample in place. In Figure 4(c), we show both experimentally measured diffracted Stokes values alongside that from our model. The fitting parameters provide good agreement between our model at normal incidence. These parameters show significant phase retardance. We observe small deviations in $S_{2}$ between our model and experiment, which may be the result of the misalignment of the incident angle of the sample or variations in PANI surface flatness.
Figure 5: CP detection with PANI-HS encoding and LP filters. (a) Experimental setup where RCP and LCP light ($P_{in}$) were generated using a polarizer and waveplates. The polarization of the scattered light ($P_{out}$) was measured by a camera assembled on a moving rail with four LP detectors (blue: 90 LP, yellow: -45 LP, purple: +45 LP, and black: 0 LP). (b) Camera images and vertically integrated camera traces. Notice that the purple (+45 LP) and orange (-45 LP) traces are reversed for RCP and LCP.
PANI-HS’s diffraction-dependent polarization transformations are similar to the angular optical transformations generally associated with the nonlocal effects of metasurfaces or metagratings 23. It can be stressed that the polarimetric sensing is possible because the azimuthal position in the Debye ring acts as the analog of a polarimeter-based rotated retarder with diattenuation in front of the detector. This is also what the Mueller matrix diffractometry images show.
@@
When light diffracts at $\phi_{0}$, the presence of weak lattice modes retard the phase of the light associated with that mode, which we associate with $\gamma_{\rm ret}$. At normal incidence, the fitting parameters to Eq. 5 are $a$ = 0.75, $\gamma_{\rm ret}$ = -70, and $\gamma_{\rm rel}$ = 85. The fact that there is close agreement between our model and measurements indicates that the eigenmodes are linearly-polarized parallel and perpendicular to the $k$-vector at angle $\phi_{0}$. These parameters indicate that the polarization modulation is distinctly measurable.
@@
We measure that the far-field degree of polarization is over 95%; for our thin monolayer encoder, the assumption of polarization coherence is reasonable. Full transformations from linear-to-circular polarization are observed with non-normal angles of incidence when light is composed of both TE and TM polarizations (Supporting information Fig. S4). The full transformation from linear-to-circular polarization is observed with non-normal angles of incidence when light is composed of both TE and TM polarizations at an 8.5-degree angle of incidence. With the ordering ratio of $\kappa/\kappa_{0}\leq 1.5$, we observe circular-to-linear polarization conversion as well as linear-to-circular, and note that other polarizations are tuned with an angle of incidence in a manner analogous to metasurfaces that enable full-Stokes transformation and polarization tuning.
[…]
To sense the handedness of the CP light, the key asymmetry arises in the tilted axes of the diffracted polarization; if this retardation is not present, then the CP component is not sensed. This effective birefringence is intriguing. The tilt of the linear polarization axis, also known as the azimuth of the polarization ellipse, depends on the CP handedness. The diffracted light appears as if it passes through a rotated retarder and diattenuator; when PANI-HS is illuminated with RCP and LCP, the Debye ring appears as a vector vortex ring, where the azimuth of the polarization ellipse is opposite for RCP and LCP. With an LP camera (Thorlabs CS505MUP1), four similar images with 0, 45, 90, and 135-angled LP filters are taken and summed vertically (along the axis perpendicular to the diffraction) [Fig. 5(a)]. As shown in Fig. 5(b), RCP and LCP are differentiated by the +45 and -45-LP components, which flip in power depending on the circular-polarization helicity. The relative strength between these two detected +45 and -45-LP signals reveals the relative phase between transverse electric (TE) and transverse magnetic (TM)-polarization components. This encoder with an LP camera achieves full-Stokes sensitivity.
@@
We draw direct analogy to compressed sensing where there are two fundamental requirements: 1) the input information is sparse 2) the information is encoded incoherently 5. This second requirement is addressed by disorder in the sample 24. We refer to the ‘ordered’ and ‘random’ samples characterized in Fig. 2. Figures 6(a-b) illustrates the encoder wavefront distortion or incoherence function, which is stronger with random samples. When the sample is ordered (polycrystalline, or carries sharp pair-wise correlations and larger grains), then light is directed into angularly-sharp $k-$space diffraction peaks. When the sample is random and has short-range order, then the diffraction patterns are blurred. When the sample contains multiscale features such as nanofibers, the signal is spatially multiplexed into multiple $k-$space modes. The mode broadening and multiplexing enhances the capability for the single input-angle information to be captured in a few measurements.
@@
Experimentally, the output angle of the Debye ring $\theta_{o}$ follows a simple grating relation, however the peak power values at this diffracted angle varies approximately 2dB or 30%. In Fig. 6(c-d) we plot the peak transmitted power as a function of input angle captured for ordered and random structures. This reveals the interference patterns in the diffracted light patterns, which are more evident for ordered than random samples. The interference fringes are regularly spaced when we plot the peak intensity as a function of $k$ or $\cos(\theta_{o})$. The fringes are easily interpreted as noise: one must sample the input angle $\theta_{i}$ with high resolution in order to observe the fringes clearly. These sharp interference fringes, which range in angular spacing $3-8^{\circ}$ from $\theta_{i}=-75$ to $+75^{\circ}$, encode beam pointing.
@@
With simulations of ideal periodic lattices with effective medium theory (i.e., low refractive indices due to the air-filled fraction of the spheres), we attribute the polarization modulation to the presence of Rayleigh anomalies or TE-polarized lattice modes at $\theta_{i}=\pm 20^{\circ}$. Experimentally, this angle coincides with increased absorption and diffraction and strong polarization modulation. The ${\rm S}_{2}$ Stokes measurements or the tilt of the azimuth of the polarization ellipse, varies significantly between RCP and LCP inputs [Fig. 6 (e-f)]. To underline why the hollow-sphere structure is important: simulations do not show the presence of the Rayleigh anomaly with solid PS (with a higher refractive index) in similar lattice geometries. These results are consistent with prior experimental observations that lattices composed of PS spheres of similar dimensions do not significantly transform the diffracted light polarization like PANI-HS.

### v2

@@ Abstract
Sensing light’s polarization and wavefront direction enables surface curvature assessment, material identification, shadow differentiation, and improved image quality in turbid environments. Traditional polarization cameras utilize multiple sensor measurements per pixel and polarization-filtering optics, which result in reduced image resolution. We propose a nanophotonic pipeline that enables compressive sensing and reduces the sampling requirements with a low-refractive-index, self-assembled optical encoder. These nanostructures scatter light into lattice modes, which encode the wavefront direction and the polarization ellipticity in the linearly-polarized components of the diffracted, interference patterns. Combining optical encoders with a neural network, the system predicts pointing and polarization when the interference patterns are adequately sampled. A comparison of “ordered” and “random” optical encoders shows that the latter both blurs the interference patterns and achieves higher resolution. Our work centers on the unexpected modulation and spatial multiplexing of incident light polarization by self-assembled hollow nanocavity arrays as a class of materials distinct from traditional metasurfaces that will not only enable encoding for polarization and optical computing but also for compressed sensing and imaging.
@@
Compressed sensing, polarimetric sensing, hollow nanospheres, colloidal self-assembly, photonic crystal thin films.
@@ 1 Introduction
The capacity to sense both light polarization and its wavefront or pointing direction offers the ability to infer the surface curvature, material, or texture, distinguish shadows from objects, see otherwise-transparent shapes, or improve image quality in turbid environments 1, 2, 3, 4. However, as independent qualities of light, polarization and pointing are typically detected separately with arrays of sensors and polarization-filtering optics. Polarization states are characterized by the Jones or Stokes parameters, i.e., four separate readings that identify the power (${\rm S}_{0}$), 2-$D$ linear-polarization (LP) axis (${\rm S}_{1}$, ${\rm S}_{2}$), and circular-polarization (CP) ellipticity (${\rm S}_{3}$). Since each polarization-filtered sensor provides a projection of the incoming wave, multiple sensor measurements are needed for each sampled pixel; “full-Stokes” cameras involve inevitable reductions in image resolution.
The process of physical encoding would mitigate the challenge of reduced resolution caused by polarization sensitivity through compressive sensing. Compressive sensing is a revolutionary technique that exploits the sparsity of a signal so that the data acquisition and storage requirements are significantly reduced without sacrificing the reconstruction quality 5. By contrast, conventional imaging is a relatively cumbersome direct-sampling approach [Fig. 1(a)]: it involves rasterized or serial measurements to identify pointing and polarization with polarization filters and multiple-pixel sensors. With conventional imaging, the grid sampling dimension $\Lambda$ must be less than the diameter of the diffraction-limited spot d, or $\Lambda<\textit{d}$. Moreover, the polarization inferred is generally only associated with the polarization that is measured. A physical encoder that aids compressed sensing maps the input optical fields over different output field components in a manner analogous to and generally paired with computing algorithms [Fig. 1(b)].
@@
where the sensing matrix $\mathbf{A}$ transfers the vector of input measurements $\mathbf{x}$ to a $1\times 4$ sensor output $\mathbf{y}$: this sensing approach is inefficient because many measurements are needed and there are many zero-valued elements in the sensing matrix $\mathbf{A}$. These zero-valued elements signify that the algorithmic operations performed by $\mathbf{A}$ are inefficient and yield no information.
Figure 1: Basic schematic of conventional polarimetric imaging and proposed polarimetric compressed sensing. For a diffraction-limited spot with diameter d, the Nyquist sampling limit refers to the minimum sampling dimension needed to capture the spot size or a grid sample dimension of $\Lambda<\textit{d}$. (a) With a conventional polarimetric camera, the direction and polarization of a point source are measured with independent and separate filters and pixels. Each image vector $\mathbf{x}$ contains many zero-valued readings. (b) With encoded and hybrid sensing, the point light source and its polarization are mapped to multiple pixels with different polarizations. Each encoded image vector $\mathbf{\alpha}$ carries joint statistics of the pointing and polarization.
Figure 2: Characterization of the optical encoders. Digital image (a) and SEM image (b) of the PANI-HS optical encoder. Scale bar is 500 nm. (c) Cartoon illustration showing the cross section of the PANI-HS optical encoder. (d) Grain boundary analyses for “ordered” (top) and “random” samples (bottom) showing the grain axes rotation (degrees). Microscope images are 30 $\mu$m x 22 $\mu$m. (e) The pair correlation function $g(r)$ computed from optical images against the normalized radius $r/2r_{0}$, where $r$ is the distance from an arbitrary origin, and $r_{0}$ is the mean sphere radius. The full width at half maximum (FWHM) of the first peak in the Fourier transform of the function $g(r)^{-1}$ provides a measure of the structural disorder. Perfect lattice (top), example “ordered”(middle) and “random” sample (bottom) are shown.
@@
Here, we demonstrate the compressed sensing of polarized signals with low-refractive-index, self-assembled encoders. The unexpected transformations from linear to elliptical polarizations in the self-assembled grating structures have only recently been reported in our work 7. We draw an analogy to the compound-lens imaging system of insects, which is covered with self-assembled corneal random and ordered grating nanostructures 8, 9, 10. Although nanostructures in nature are associated with anti-reflective, hydrophobic, or self-cleaning functions 11, 12, the optical encoding function of their nanostructures has not been explicitly shown. A survey of the potential modes for insect-inspired imaging and object tracking (colorimetric, polarimetric filtering, coherent optical preprocessing, and compressive sensing) has recently been written 13. Whether insect corneal nanostructures have a physical encoding function or not, our approach with self-assembled encoders illustrates the potential.
A key component to our encoders is that they scatter light into lattice modes that are sensitive to the direction of the incoming wave and its polarization ellipticity. Our work not only expands on our understanding of the physical encoding that is available with self-assembled nanostructures, but also outlines general insight for their design with compressed sensing.
@@ 2 Results
Figure 3: Polarization transformations of PANI-HS compared with those of polystyrene spheres (PS). (a) (i-ii) SEM images of PS and PANI-HS (720-nm lattice spacing) and corresponding (iii-iv) Debye or diffracted ring intensity patterns from vertical, linearly polarized (LP) light (532-nm wavelength). The measured diffracted polarization is overlaid in yellow. For PS, the diffracted intensity is uniform and the diffracted polarization remains vertical but for PANI-HS, both intensity and polarization vary spatially and the polarization ellipticity indicates spatially varying birefringence. (b) Conoscopy measurements showing (i-ii) diattenuation and (iii-iv) retardance magnitude and phase for single-crystalline PANI-HS, polycrystalline PANI-HS, and PS with 2-$\mu$m lattice spacing. Retardance and diattenuation are negligible in the PS samples compared to those of PANI-HS.
@@
We fabricate hierarchical 2D photonic-crystal films composed of hydrochloric acid-doped polyaniline nanofibers around hollow spheres (PANI-HS) based on an approach demonstrated in our prior work 7. Figures 2(a-c) show the air-filled polymer grating structures. The film structure at the smallest level is composed of nanofibers surrounding a sub-micron-scaled spherical cavity, shown by the SEM image in Fig. 2(b). We prepare PANI-HS films with two types of structures: polycrystalline (“ordered”) and short-range ordered (“random”) to investigate the impact of ordering on compressive sensing. Structural analysis was conducted using boundary counts and the 2D pair correlation function.
Figure 2(d) shows the orientation of different grains, revealing boundary counts of 1.8 boundaries/$\mu$m and 3.3 boundaries/$\mu$m for the ordered and random samples, respectively. In the boundary count method 53, a grain is defined within a lattice with specific orientation; for our 6-fold symmetric assembly, a grain’s orientation relative to an axis ranges from 0 to 60 degrees. A grain boundary occurs when neighboring domains deviate by a small angle. Fig. 2(e) shows the 2D pair correlation function $g(r)$ 14, 15, 16 of the ordered and random samples alongside a perfect lattice. We also assessed ordering by comparing the width of the first-order peaks in the Fourier transform to that of a perfect lattice, yielding average orderings of 1.10 and 2.49 for ordered and random samples across three randomly selected areas. Additional detailed structural analysis is provided in Supporting Information Fabrication and Material Characterization of PANI-HS (Figs. S1 and S2).
[…]
Figure 4: Polarization transformations with normally incident light. (a) Experimental setup where RCP and LCP light ($P_{\rm in}$) are produced with polarizer and waveplates (WP). Images at different polar angles $\theta$ are taken with a linear-polarization (LP) camera. (b) Linearly polarized images ($P_{\rm out}$) and vertical line traces over each polar angle (blue: 90LP, yellow: -45LP, purple: +45LP, and black: 0LP). Notice that the purple (+45 LP) and orange (-45 LP) traces are reversed for orthogonal input circular polarizations ($P_{\rm in}$). Modeled and measured Stokes parameters (c) for the input, where -90 (TE), 0 (TM), and 90 (TE) are the polarization angles with respect to the plane of measurement and (d) at the diffracted, Debye-ring polar angle ($\theta=56^{\circ}$).
@@
To illustrate the polarization transformations of the film, we use Mueller matrix diffractometry, which shows the angle-dependent diattenuation (or polarization-sensitive losses) and retardance of the PANI-HS and PS arrays 19 [Fig. 3(b)]. In these measurements, the Mueller matrix polarimeter images the rear focal plane of the 50x (0.95 NA) objective to obtain complete polarization-resolved maps of the directional distribution of light. The complete Mueller matrix for each image in Fig. 3(b) is provided in the Supporting Information Diffractometry Measurements (Fig. S3). The high magnitudes of diattenuation and retardance in PANI-HS indicate that the material exhibits strong polarization-selective properties and significant birefringence, which is less pronounced in the PS film. The comparison of measurements that probe single crystals vs. grain boundaries across multiple crystals [Figs. 3(b)i-ii] shows similar diattenuation and retardance patterns. This symmetry of the diattenuation and retardance allows us to rule out stress-induced birefringence in the coating 20. Also, since similar trends are observed for single and multiple crystals, the birefringence is not associated with the grain boundaries or defects in the self-assembled films 21.
The PANI-HS response to incident polarization ellipticity is measurable with an LP camera or a camera with LP-filtered sensor pixels [Fig. 4(a-b)]. To sense the handedness of the CP light, the key asymmetry arises in the tilted axes of the diffracted polarization; if this retardation is not present, then the CP component is not sensed. This effective birefringence is intriguing. The tilt of the linear polarization axis, also known as the azimuth of the polarization ellipse, depends on the CP handedness. The diffracted light appears as if it passes through a rotated retarder and diattenuator; when PANI-HS is illuminated with RCP and LCP, the Debye ring appears as a vector vortex ring, where the azimuth of the polarization ellipse is opposite for RCP and LCP. With an LP camera (Thorlabs CS505MUP1), four similar images with 0, 45, 90, and 135-angled LP filters are taken and summed vertically (along the axis perpendicular to the diffraction) [Fig. 4(a)]. As shown in Fig. 4(b), RCP and LCP are differentiated by the +45 and -45-LP components, which flip in power depending on the circular-polarization helicity. The relative strength between these two detected +45 and -45-LP signals reveals the relative phase between transverse electric (TE) and transverse magnetic (TM)-polarization components. This encoder with an LP camera achieves full-Stokes sensitivity.
@@
The diffraction from nanofibrous PANI-HS films appears as the “pooled” scattered light from a single dipole scattering event with additional birefringence. A simple, coherent model fit to our measurements identifies scattering and retardance as the two underlying contributions to the polarization transformation. The polarization transformation is viewed as a rotated polarizer $\mathbf{J}_{\rm pol}$ and rotated waveplate $\mathbf{J}_{\rm wp}$,
@@
where $\phi_{0}$ is measured in the sample plane with respect to the $x$-axis, and $\mathbf{R}(\phi_{0})=\begin{bmatrix}\cos(\phi_{0})&-\sin(\phi_{0})\\ \sin(\phi_{0})&\cos(\phi_{0})\end{bmatrix}$ is the rotation matrix associated with azimuthal angle $\phi_{0}$, which is the LP polarization angle with respect to the incident plane.
@@
At normal incidence, the fitting parameters to Eq. (5) are $a$ = 0.75, $\gamma_{\rm ret}$ = -70, and $\gamma_{\rm rel}$ = 85. These parameters show significant retardance. Figures 4(c-d) show the input and transformed Stokes parameters generated by this model. The close agreement between our model and measurements indicates that the eigenmodes are linearly-polarized parallel and perpendicular to the $k$-vector at azimuthal angle $\phi_{0}$. These parameters indicate that the polarization modulation is distinctly measurable. The small deviations in ${\rm S}_{2}$ between our model and experiment may be attributed to misalignment of the incident angle of the sample or variations in the PANI surface flatness.
Figure 5: Numerically simulated Jones polarization plots of the diffracted mode by the PANI-HS (second row) and PS solid spheres (third row) arranged in a hexagonal lattice. The lattice orientation and incident linear polarization are shown by red arrows (first row).
It can be stressed that the polarimetric sensing is possible because the azimuthal position in the Debye ring acts as the analog of a polarimeter-based rotated retarder with diattenuation in front of the detector. This is also what the Mueller matrix diffractometry images show (See Supporting Information).
@@
We measure that the degree of polarization in the far field is over 95%; for our thin monolayer encoder, the assumption of polarization coherence is reasonable. The full transformation from linear-to-circular polarization is observed with non-normal angles of incidence when light is composed of both TE and TM polarizations at an 8.5-degree angle of incidence (Supporting Information Fig. S4). We observe circular-to-linear polarization conversion as well as linear-to-circular, and note that other polarizations are tuned with an angle of incidence in a manner analogous to metasurfaces that enable full-Stokes transformation and polarization tuning.
To further validate the developed model (see Eqs. 5-7) and gain additional physical insights on the polarimetric transformation of PANI-HS, we perform full-wave numerical analyses. Using the commercial software Comsol Multiphysics, we simulate the PANI-HS and PS arranged in an ordered hexagonal lattice. Details on the numerical simulations are reported in the Methods section. Figure 14 shows the polarization of the diffracted mode describing the Debye ring for four different orientations of the electric field of the incoming LP plane wave at normal incidence. When the electric field of the incident wave is parallel to the horizontal and vertical axis of the lattice (see row 1 and columns 2 and 5 in Fig. 14), neither PANI-HS nor PS structures transform the polarization and the diffracted mode remains LP, similar to the incident wave.
On the other hand, for an incident LP wave with an electric field not directed along either the horizontal and vertical axes of the lattice (see row 1 and columns 3 and 4 in Figs. 14), the PANI-HS structure performs a polarization transformation. The PANI-HS diffracts an elliptically polarized mode, while the PS structure diffracts mostly LP light. These results show excellent agreement with the experimental results [Fig. 3(a)], confirming the ability of the PANI-HS to manipulate light polarization. They also show the accuracy of the developed model [Fig. 4(c-d)], which predicts the linear-to-elliptical polarization conversion of the PANI-HS. When the direction of the electric field of the incident wave deviates from the horizontal axis, the amplitude of the Stokes parameter S3, which indicates the strength of elliptically polarized light with S1 is different than zero, increases and returns to zero for a vertically polarized incoming wave.
@@
PANI-HS’s diffraction-dependent polarization transformations are similar to the optical transformations generally associated with the nonlocal effects of metasurfaces or metagratings 22. Notably however, an important aspect of the self-assembled encoder is a non-negligible degree of disorder, which offers an “incoherence” function for compressive sensing. Compressed sensing is defined by two fundamental requirements: 1) sparse input information (in the conventional imaging system [Fig. 1(a)]), and 2) “incoherent” output information 5 (with the PANI-HS encoded system [Fig. 1(b)]). This second requirement is not only influenced by the PANI-HS geometry but also addressed by disorder in the sample 23.
@@
We refer to the “ordered”and “random” samples characterized in Fig. 2. The random PANI-HS samples offer a higher degree of incoherence compared to the ordered ones with broadened $k$-space angular modes, less-distinct fringes, and broadened modulation of the far-field polarization. Figures 6(a-b) illustrate the encoder wavefront distortion or incoherence function. When the sample is ordered (polycrystalline, with sharp pair-wise correlations and larger grains), then light is directed into angularly-sharp $k-$space diffraction peaks. When the sample is random and has short-range order, then the diffracted interference patterns are blurred. When the sample contains multiscale features such as nanofibers, the signal is spatially multiplexed into multiple $k-$space modes. The mode broadening and multiplexing enhances the capability for the single input-angle information to be captured in a few measurements.
@@
Experimentally, the output angle of the Debye ring $\theta_{o}$ follows a simple grating relation, however the peak power values at this diffracted angle varies approximately 2dB or 30%. In Figs. 6(c-d) we plot the peak transmitted power as a function of input angle captured for ordered and random structures. This reveals the interference patterns in the diffracted light patterns, which are more evident for ordered than random samples. The interference fringes are regularly spaced when we plot the peak intensity as a function of $k$ or $\cos(\theta_{o})$. The fringes are easily interpreted as noise: one must sample the input angle $\theta_{i}$ with high resolution in order to observe the fringes clearly. These sharp interference fringes, which range in angular spacing $3-8^{\circ}$ from $\theta_{i}=-75$ to $+75^{\circ}$, encode beam pointing.
@@
The diffracted polarization also varies as a function of input angle for ordered and random structures. This can be explained by the presence of weak lattice modes, which retard light and depend on the light’s incident input angle and polarization. With simulations of ideal periodic lattices with effective medium theory (i.e., low refractive indices due to the air-filled fraction of the spheres), we attribute the polarization modulation to the presence of Rayleigh anomalies or TE-polarized lattice modes at $\theta_{i}=\pm 20^{\circ}$ (See Supporting Information). The refractive index data from prior PANI characterization 56 is used. Experimentally, this angle of strong polarization modulation also coincides with increased absorption and diffraction. The ${\rm S}_{2}$ Stokes measurements or the tilt of the azimuth of the polarization ellipse, varies significantly between RCP and LCP inputs [Fig. 6 (e-f)]. To underline why the hollow-sphere structure is important: simulations do not show the presence of the Rayleigh anomaly with solid PS (with a higher refractive index) in similar lattice geometries. These results are consistent with prior experimental observations that lattices composed of PS spheres of similar dimensions do not significantly transform the diffracted light polarization like PANI-HS.